The manager of an electric supply store measured the diameters of the rolls of wire in the inventory. The diameters of the rolls (in m) are listed below. Calculate the standard deviation.

0.523    0.139    0.466    0.403   0.343    0.13    0.482
Answer #1

Sample mean \bar{x } = \sum X / n

= ( 0.523 + 0.139 + 0.466 + 0.403 + 0.343 + 0.13 + 0.482) / 7

= 2.486 / 7

= 0.35514

Standard deviation S = Sqrt( \sum X2 - n\bar{x }2 / n-1)

= sqrt [ ( ( 0.5232 + 0.1392 + 0.4662 + 0.4032 + 0.3432 + 0.132 + 0.4822) - 7 * 0.355142 ) / 6 ]

= 0.1615